OpenStudy (anonymous):

(x²-4)/(x²-x-6) divide and simplify the expression

OpenStudy (anonymous):

note that (x^2 - 4) = (x+2)(x-2) and (x^2 - x - 6) = (x +2)(x-3) i can answer follow up questions :)